OpenStudy (anonymous):

would it be A or C? How would the acceleration of a chain of three shopping carts compare with the acceleration of a single cart if the same force acted on both? A. 1/3 as much the single B. 1/2 as much the single C. 3 times as much the single D. 2 times as much the single

OpenStudy (anonymous):

a = F/m a' = F/3m a'/a = 1/3